OpenStudy (unklerhaukus):

True or False \[\mathbb{N}\in\mathbb{Z}\in\mathbb{Q}\in\mathbb{R}\in\mathbb{C}\]

OpenStudy (kinggeorge):

Unless I'm missing some surprising insight, this should be true.

OpenStudy (blockcolder):

The symbol's wrong, I think. It should be "subset", not "element".

OpenStudy (unklerhaukus):

How is it wrong blockcolder/

OpenStudy (unklerhaukus):

because the terms are sets not points?

OpenStudy (blockcolder):

I think so. \[\mathbb{N}\subset\mathbb{Z}\subset\mathbb{Q}\subset\mathbb{R}\subset\mathbb{C}\]

OpenStudy (unklerhaukus):

ok well that makes sense,

OpenStudy (kinggeorge):

That seems to be the surprising insight I missed.